If a gas occupies 2.0 L at 0.0 °C, what is the volume at 273 °C if the pressure does not change?
Answer #1

By Charles' law, at constant pressure

Volume of a gas (V) is directly proportional to absolute temperature (T).

Or, =

Given, V1 = 2.0 L

T1 = (0.0+273.15 ) K = 273.15 K.

T2 = (273+273.15) K = 546.15 K.

Then, V2 =

Or, V2 = (2.0 × 546.15/273.15) L

= 4.0 L ( 2 significant figures).

So, volume at 2730c is 4.0 L.